A member ABCD is subjected to point loads P1,P2,P3, and P4 calculate the force P3 necessary for equilibrium if P1, = 120 kN, P2, = 220 kN and P4, = 160 kN. Determine also the. net change in the length of the member. Take E = 200 GN/m²

1
Expert's answer
2022-02-16T06:53:01-0500

"\sum P=0"

"P_1+P_2+P_3+P_4=0"

120+220+160+P3=0

P3=-500kN

"∆l=\frac{\sigma l}{E}"

="\frac{500000*25}{25*15*200*10^9}"

=1.6667*10-7